# Information theory for data-driven model reduction in physics and biology

### v2

@@ VI Relevant variable identification in laboratory-generated and atmospheric flows
The first shows a von Kármán street which forms as water passes by a cylindrical obstacle at Reynolds number 171, with flow visualized by a dye injected at the site of the obstacle [58]. We take a background-subtracted grayscale image of the flow field as our input (Fig. 5e) and task VIB with learning a two-dimensional latent variable as above. Also here, variational IB learns oscillatory dynamics of the latent variables (Fig. 5f). We visualize the function learned by the encoder by considering gradients of the latent variables, which show the same structure as those obtained for the $x$ component of the simulated data (Fig. 5g). This is expected, as the $x$-component of the velocity field has similar glide reflection symmetry as the intensity image.
@@
Next, we apply variational IB to a von Kármán street arising due to flow around Guadalupe Island, which was imaged by a National Oceanic and Atmospheric Administration (NOAA) satellite [59] (Fig. 5h). The video consists of only 62 frames, and the von Kármán street undergoes a single oscillation. Even with this small amount of data, the variational IB neural network learns latent variables which capture this oscillation and have the expected dependence on the input variables (Fig. 5i-j). As in the first experimental example, the gradients of the encoding variables show the glide symmetry of $m_{x}$ due to the symmetry of the intensity pattern in Fig. 5d. This symmetry is less clear in the component $\frac{\partial h_{1}}{\partial I(x)}$, which is likely due to the fact that the von Kármán street is not as fully formed in this data as in our previous examples.
@@ VII Relevant variable discovery in cyanobacterial populations
We were provided with videos of 10 cyanobacteria colonies that were grown under various conditions that impact their dynamics. However, as a test of our method, we were blinded to these conditions until we had performed our analysis. The videos are sequences of fluorescent images, taken once per hour, which show the clock state of each individual bacteria visualized with a fluorescent marker EYFP driven by the kaiBC promoter. Here, we focus on collective variables which are predictive of the state of the interior of the colony and not the growth in area of the colonies. We therefore crop the images to the interiors of each colony (SI Fig. S9). This allows us to isolate the motion of individual bacteria and fluorescence oscillations (Fig. 6b).
@@
Our input to the variational IB neural network are these cropped images augmented with a time-lagged image of the same region (Fig. 6c). The purpose of this time lag is to make the dynamics Markovian: due to the oscillatory intensity field, if one observes only a single time point it is unclear whether the intensity is currently increasing or decreasing. These time lagged pairs comprise our system state, $X_{t}=\{I(\mathbf{x},t),I(\mathbf{x},t+\tau)\}$, where $\tau$ is the duration of the time lag. Here we take $\tau=3$ hr and a prediction time horizon $\Delta t=8$ hr, but find that choosing different $\Delta t$ or $\tau$ does not change our results (SI Fig. S9).
@@
Figure 6: Discovering slow collective variables in cyanobacteria populations (a) Fluorescent images of cyanobacteria colonies labeled with EYFP driven by the kaiBC promoter, allowing the visualization of Kai protein transcription. The colonies are imaged as they undergo cell growth and oscillations in Kai expression associated with the circadian rhythm. (b) Time series of one cropped section of an individual colony (see SI Fig. S9 for details). (c) The “state” used for variational IB is the time-lagged intensity field with lag time $\tau$. (d) Variational IB embeddings of time-lagged images into two dimensions (see Supplementary Movie 2). Every line corresponds to one colony’s evolution in time. Note the apparent oscillations of different radii. (e) Embedding into three dimensions. We orient our axes to correspond to the three principal components of the data. Thus, the projection onto the $h_{2}=0$ plane corresponds to a projection along the dominant two principal components. In this subspace, we see a similar structure as in the 2D embedding. (f) Mutual information between the future state and the encoding, given by $I(X_{t+\Delta t},H_{t})$, for varying dimension of the latent space. Increasing the dimension of the embedding space beyond two leads only to marginal increases in $I(X_{t+\Delta t},H_{t})$; this tells us “when to stop.” Each small point represents one training instance of the variational IB model, while the large point shows the maximum estimated value. Because the InfoNCE estimator is a lower bound on the true information, we consider the maximum as the estimated mutual information. The value of $I_{\max}=I(X_{t},X_{t+\Delta t})$ is the mutual information estimated for the true dynamics. (g) Selected trajectories in latent space with large and small radii. (h) Time-average latent radius of all colonies. (i) Mean synchronization order parameter of the intensity images; see SI for computation details. (j) Mean radius versus synchronization parameter for each colony. VIB identifies clusters of cells characterized by high (blue) or low (black) synchronization. As revealed to us after our analysis, this clustering corresponds to differing theophylline concentrations across experiments. Within each experimental movie (each of which contains 2-3 colonies) the radii are mostly constant. (k) Sample time series of weakly (black) and highly (blue) synchronized colonies. We apply a slight Gaussian blur to better visualize the bacteria boundaries.
@@
With variational IB we compress the state $X_{t}$ into a latent variable $h$ of variable dimension (Fig. 6d-f). We train the neural network on the entire dataset of all 10 colonies simultaneously. The dynamics in latent space undergo clear oscillations, indicating that the relevant variables encode primarily the intensity fluctuations rather than, for example, the spatial locations of the bacteria. Notably, the trajectories are essentially two-dimensional, even when the encoding space is higher dimensional. This is reflected in the information retained about the future state, $I(X_{t+\Delta t},H_{t})$. We see that increasing the dimension of the embedding space beyond two leads only to marginal increases in $I(X_{t+\Delta t},H_{t})$; this tells us “when to stop” (Fig. 6f). We independently verify this by using principal component analysis to characterize the geometry of embedded trajectories, and find that even in higher dimensions the trajectories occupy a two dimensional subspace (SI Fig. S10). In the following, we therefore restrict our focus to the $\dim H=2$ case.
@@
We noticed that there were notable differences in the radius of latent space oscillations from colony to colony, two of which are highlighted in (Fig. 6g). To understand this difference, we examined the original microscopy time series corresponding to both large and small latent radius (Fig. 6k) and found that while the large-radius sample showed clear, nearly uniform oscillations in intensity, the small-radius samples appeared much more heterogeneous.
@@
To quantify this we consider each pixel to be an independent oscillator, akin to a spatial Kuramoto model [62, 63, 64], and compute a global synchronization order parameter $r(t)$ (see SI). For each colony we calculate the time-averaged synchronization $\langle r(t)\rangle_{t}$ and find that two clusters emerge corresponding to high and low synchronization (Fig. 6i). These clusters are precisely those representing trajectories of large and small latent radius (Fig. 6j), suggesting that variational IB learns to encode the synchronization of the colony in the latent variable radius. As a check, we perform IB on a simulated locally-coupled Kuramoto model as a system which shares many features of the experimental system. Here we also learn an encoding in which the latent radius corresponds to the synchronization order parameter (SI Fig. S11).
@@
In the SI we compare the performance of variational IB to several other model reduction methods and find that IB delivers more interpretable and well-behaved features. This is likely due to the fact that many standard methods for data-driven model reduction rely on assumptions about the dynamics which may not be appropriate in the case at hand, such as linearity. Even among deep learning methods such as time-lagged autoencoders that are free of such assumptions, the variables learned by IB appear more interpretable. This increased interpretability is likely due to the compression term which effectively regularizes the latent space by encouraging the network to learn slow transfer operator eigenfunctions. While there are many specific use variants of DMD [13, 65, 66, 67, 68] or autoencoders for dynamics [20, 21, 22, 69] that may outperform variational IB in some cases, we find that in this real-world example it yields the smoothest and most interpretable latent variables without requiring tailored pre-processing steps (SI Fig. S12).
@@ VIII Conclusion
We have related information-theoretical properties of dynamical systems to the spectrum of the transfer operator. We illustrate our findings on several simple and analytically tractable systems, and turn them into a practical tool using variational IB, which learns an encoding variable with a neural network. The latent variables of these networks can be interpreted as transfer operator eigenfunctions even though the network was not explicitly constructed to learn these: it optimizes a purely information-theoretic objective that contains no knowledge of a transfer operator or dynamics. This allows one to harness the power of neural networks to learn physically-relevant latent variables. Biological systems are an ideal setting for such methods: despite their apparent complexity, they can often be captured by low-dimensional descriptions which are difficult to identify by physical considerations alone [70, 47, 71, 72]. We have shown that variational IB is a potentially powerful tool for these cases, and can discover slow variables even directly from image data without significant preprocessing.
Methods
Mutual information and entropy
$S(X)=-\int\text{d}x\,p(x)\log p(x).$
Given another random variable $Y$, such that $X$ and $Y$ have a joint distribution $p(x,y)$, we can ask how much information is shared between these two variables. This is given by the mutual information
$I(X,Y)$ $=S(X)-S(X|Y)$ (12)
$=D_{\text{KL}}(p(x,y)\lVert p(x)p(y))$ (13)
This can be interpreted as quantifying how much (on average) a measurement of $Y$ can reduce our uncertainty about the value of $X$ ((12)).
@@ Methods
We are concerned with the sign of the eigenvalues of $H$. A negative eigenvalue indicates that $\mathcal{L}_{\text{IB}}$ is unstable to a perturbation in $f$, which means the loss can be lowered by changing $f$ away from the trivial encoder at $f_{n}=0$. Because the eigenvalues of a tensor product of matrices are products of the eigenvalues of the component matrices, the eigenvalues of $H$ change sign when those of $A^{\beta}$ do. $A^{\beta}$ and its spectrum can be computed, which we do in the SI. The result of this calculation is that the first eigenvalue to become negative is associated with the eigenvector $\mathbf{v}=(1,0,0,...)$. This computation is exact for equilibrium systems, which are those in which the steady-state flux vanishes, but in nonequilibrium systems there may generally be a correction proportional to the flux. In summary, this means that in the limit of high-compression only the first component $f_{1}$ becomes non-zero, hence the encoder has the form given in (19).
@@ Acknowledgements
Appendix A Information bottleneck
@@
$\mathcal{L}_{\text{IB}}[p(h|x)]$ $=I(X,H)-\beta I(Y,H).$ (20)
@@
$\begin{split}\mathcal{L}_{\text{IB}}[p(h|x)]&=\sum_{x,h}p(h|x)p(x)\log\frac{p(h|x)}{p(h)}-\beta\sum_{y,h}p(y,h)\log\frac{p(y,h)}{p(h)p(y)}+\sum_{x}\lambda(x)\left(1-\sum_{h}p(h|x)\right).\end{split}$ (21)
@@
When $\beta<1$, it follows from the data processing inequality $I(X,H)\geq I(Y,H)$ that (20) is minimized by a trivial encoder $p(h|x)=p(h)$. In this case, $\mathcal{L}_{\text{IB}}=0$ and no information passes through the bottleneck. As $\beta$ is increased, more information is allowed through the bottleneck until the relevant variables begin to gain a dependence on the state $x$. This occurs suddenly for a certain value of $\beta=\beta_{1}>1$ at which the encoder becomes non-uniform and $I(H,X)$ becomes non-zero. This is referred to as an IB transition; for increasing $\beta$, there may be a sequence of transitions at $\beta_{2},\beta_{3},...$ etc.
Appendix B Numerically solving “exact” IB and Ulam approximations of the transfer operator
The optimal IB encoder (22) can be found using the Blahut-Arimoto (BA) algorithm [43]. As described in detail in Refs. [30, 79] and sketched in Fig. S1, the algorithm is an iterative procedure, where the encoder at iteration $k+1$ is updated according to
The first line simply plugs the previous estimate of the encoder in to Eq. 29 (and normalizes the distribution), while the following two lines update marginal and conditional distributions using the new estimate of the encoder. In Refs. [30, 79] it is shown that this algorithm converges.
$P_{ij}=P(X_{t+\Delta t}=x_{j}|X_{t}=x_{i})=N_{i\rightarrow j}/N_{i},$ (24)
$Up(x_{j})\approx\sum_{x_{i}}P_{ij}p(x_{i}),$ (25)
Appendix C Transfer operators
@@
$\dot{\mathbf{x}}=\mathbf{f}(\mathbf{x})+\sqrt{2D(\mathbf{x})}\mathbf{\xi}(t)$ (26)
@@
This framework captures purely deterministic dynamics, which are obtained by setting $D=0$. The transfer operator describes the evolution of probability distributions. For dynamics given by a Langevin equation, probability distributions evolve according to the corresponding Fokker-Planck equation
@@
Here we recognize $\mathcal{L}$ as the infinitesimal generator of the transfer operator $U$: evolution for a time $t$ is given by $U=\text{e}^{t\mathcal{L}}$.
The adjoint operator is given by the so-called backward Kolmogorov equation, and it describes the evolution of functions $\phi$.
$\mathcal{L}^{\dagger}\phi(x)=f_{i}(\mathbf{x})\partial_{i}(\phi(\mathbf{x}))+D(\mathbf{x})C_{ij}\partial_{i}\partial_{j}(\phi(\mathbf{x})).$ (27)

### v3

@@ V Data-driven discovery of slow variables
Our framework provides interpretability for the learned latent variables even in messy real-world fluid flow datasets scraped directly from videos on Youtube [73, 75] (Supplementary Movie 1). The first shows a von Kármán street which forms as water passes by a cylindrical obstacle at Reynolds number 171, with flow visualized by a dye injected at the site of the obstacle [73]. We take a background-subtracted grayscale image of the flow field as our input (Fig. 5d) and task VIB with learning a two-dimensional latent variable as above. Also here, variational IB learns oscillatory dynamics of the latent variables (Fig. 5e). We visualize the function learned by the encoder by considering gradients of the latent variables, which show the same structure as those obtained for the $x$ component of the simulated data (Fig. 5f). This is expected, as the $x$-component of the velocity field has similar glide reflection symmetry as the intensity image.
@@
We also apply variational IB to a von Kármán street arising due to air flow around Guadalupe Island, which was imaged by a National Oceanic and Atmospheric Administration (NOAA) satellite [75] (Fig. 5g). Although the video consists of only 62 frames and less than one full oscillation, the variational IB neural network learns latent variables which capture this oscillation and have the expected dependence on the input variables (Fig. 5h-i). As in the first experimental example, the gradients of the encoding variables show the glide symmetry of $m_{x}$ due to the symmetry of the intensity pattern in Fig. 5g.
Figure 6: IB as part of an equation learning pipeline. (a) Kymographs of the sine-Gordon system, discussed in the text. Left shows the field $\phi(x)$ while the right shows time-derivative of the field. Initial conditions are shown above as curves. (b) Information content contained in the encoding for varying dimensionality of the embedding space. While the information appears to plateau at $D^{*}=2$, this does not necessarily mean a “proper” embedding dimension has been reached. (c) Estimated fractal dimension of the embedded dynamics for varying embedding dimension. The full fractal dimension of $\approx 2$ is only attained with an encoding dimension $D^{*}=3$. (d) Error of the predictions for the latent variable evolution produced by the dynamical system learned by SINDy. The error reaches a minimum at $D^{*}=4$. Interestingly the error begins to rise for $d>4$, likely due to the fact that the IB encoding begins to store other features of the full state variable which are irrelevant for the dynamics. (e) For $d=2$, the encoding manages to project the full torus into two dimensions by cutting and unfolding it (left). This introduces discontinuous jumps in the evolution of $h$. (f) Time evolution of both components of $h$, with highlighted part corresponding to the trajectory highlighted in (e). (g) Trajectory for a four-dimensional latent variable, visualized by projected onto the first three principal components. True trajectory shown in black, trajectory predicted by SINDy shown in blue with the equations given in the main text. (h) Predicted trajectory for all the degrees of freedom of $h$.
VI Learning the dynamics of latent variables
In Section I we showed how using an information-theoretic measure of predictability reduces the problem of model reduction to the problem of identifying relevant variables. In this setting, the dynamical rule governing the evolution of $H_{t}$ is not explicitly learned, but is instead induced by the choice of encoding variable. We now show how variational IB can be incorporated into a complete model reduction pipeline by combining it with sparse equation learning methods such as SINDy [76, 77]. We consider a system described by an underdamped, driven, one-dimensional PDE known as the sine-Gordon equation [78] (Fig. 6a; see SI Section 9). For a particular choice of driving frequency and strength, this system undergoes quasiperiodic dynamics, and hence lives on a toroidal manifold.
To identify the reduced variables for the system, we must make a choice about the dimensionality of $H_{t}$. As discussed above, the mutual information between $H_{t}$ and $X_{t+\Delta t}$ provides an indicator of “when to stop” (Fig. 6b). This metric plateaus already at $\dim H_{t}=2$. The fractal dimension of the latent manifold, however, does not plateau until $\dim H_{t}=3$ (Fig. 6c). Interestingly, when using the encoding to learn a deterministic equation for the evolution of $H_{t}$ using SINDy, the prediction error is only minimized at $\dim H_{t}=4$ (Fig. 6d).
Each of these indicators of “when to stop” tells us something different about the system. The encoding learned for $\dim H_{t}=2$ is shown in Figure 6e, with a portion of the trajectory highlighted. Although the full dynamics takes place on a torus, the encoding neural network is able to “cut” the torus to embed it in two dimensions. The dynamics of the latent variables are completely deterministic: at any point on the latent manifold it is known where the system will evolve next, so that the information about the future is complete. However, the dynamics will be highly irregular, featuring large jumps in latent space (Fig. 6f).
The “fraying” evident at the boundary of the torus embedded in two dimensions leads to a fractal dimension which is less than two, although this difference may disappear in the limit of infinite data. At three dimensions, the latent manifold is two dimensions everywhere (Fig. 6c). However, while the torus may be embedded in three dimensions, it does not admit simple, linear dynamics until $\dim H_{t}=4$. In this dimension, SINDy is able to identify the correct linear equations of motion
$\dot{h}_{0}$ $=-\omega_{1}h_{1}$ $\,\dot{h}_{2}$ $=\omega_{2}h_{3}$
$\dot{h}_{1}$ $=\omega_{1}h_{0}$ $\,\dot{h}_{3}$ $=-\omega_{2}h_{2}$ (11)
with $\omega_{1}=1.5$ and $\omega_{2}=8.7$(Fig. 6g-h).
The above example illustrates the subtle difference between a model being the simplest in the sense of providing a minimal predictive set of variables, and the simplest in terms of the governing equations. The latter may benefit from the inclusion of informationally redundant variables. The IB objective, being purely information based, is indifferent to the geometric or analytic properties of the encodings. Such model properties, often desirable, as e.g. smoothness or linearity must be introduced as additional constraints.
@@ VII Relevant variable discovery in cyanobacterial populations
We were provided with videos of 10 cyanobacteria colonies that were grown under various conditions that impact their dynamics. However, as a test of our method, we were blinded to these conditions until we had performed our analysis. The videos are sequences of fluorescent images, taken once per hour, which show the clock state of each individual bacteria visualized with a fluorescent marker EYFP driven by the kaiBC promoter. Here, we focus on collective variables which are predictive of the state of the interior of the colony and not the growth in area of the colonies. We therefore crop the images to the interiors of each colony (SI Fig. S13). This allows us to isolate the motion of individual bacteria and fluorescence oscillations.
@@
Our input to the variational IB neural network are these cropped images augmented with a time-lagged image of the same region (Fig. 7b-c). The purpose of this time lag is to make the dynamics Markovian: the intensity field oscillating, the observation at a single time point does not allow to determine whether the intensity is increasing or decreasing. In principle more time-lags may be necessary, and their number could be estimated from the behavior of the mutual information (see SI Section 3) or the transition entropy [63] . As these quantities are difficult to measure from images with limited computational resources, we keep only one lag: the state is given by pairs $X_{t}=\{I(\mathbf{x},t),I(\mathbf{x},t+\tau)\}$, where $\tau$ is the lag duration. Here we take $\tau=3$ hr and a prediction horizon $\Delta t=8$ hr, but find that different $\Delta t$ or $\tau$ does not change our results (SI Fig. S13).
@@
Figure 7: Discovering slow collective variables in cyanobacteria populations (a) Fluorescent images of cyanobacteria colonies labeled with EYFP driven by the kaiBC promoter, allowing the visualization of Kai protein transcription. The colonies are imaged as they undergo cell growth and oscillations in Kai expression associated with the circadian rhythm. (b) The “state” used for variational IB is the time-lagged intensity field with lag time $\tau$ (see SI Fig. S13 for details). (c) Variational IB embeddings of time-lagged images into two dimensions (see Supplementary Movie 2). Every line corresponds to one colony’s evolution in time. Note the apparent oscillations of different radii. Three-dimensional embedding is shown in SI Fig. S15. One large-radius (blue) and one small-radius (black) trajectory are highlighted. (d) Mutual information between the future state and the encoding, given by $I(X_{t+\Delta t},H_{t})$, for varying dimension of the latent space. Each small point represents one training instance of the variational IB model, while the large point shows the maximum estimated value. $I_{\max}=I(X_{t},X_{t+\Delta t})$ is the mutual information estimated for the full dynamics. (e) Mean radius versus synchronization parameter for each colony (see SI Section 8). VIB identifies clusters of cells characterized by high (blue) or low (black) synchronization. This clustering corresponds to differing theophylline concentrations across experiments. Within each experimental movie (each of which contains 2-3 colonies) the radii are mostly constant. (f) Sample time series of weakly (black) and highly (blue) synchronized colonies. We apply a slight Gaussian blur to better visualize the bacteria boundaries.
@@
With variational IB we compress the state $X_{t}$ into a latent variable $h$ of variable dimension. We train the neural network on the entire dataset of all 10 colonies simultaneously. The dynamics in latent space undergo clear oscillations, indicating that the relevant variables encode primarily the intensity fluctuations rather than, for example, the spatial locations of the bacteria. Notably, the trajectories are essentially two-dimensional, even when the encoding space is higher dimensional. This is reflected in the information retained about the future state, $I(X_{t+\Delta t},H_{t})$: increasing the dimension of the embedding space beyond two only marginally increases $I(X_{t+\Delta t},H_{t})$; this tells us “when to stop” (Fig. 7d). We independently verify this by using principal component analysis to characterize the geometry of embedded trajectories, and find that even in higher dimensions the trajectories occupy a two dimensional subspace (SI Fig. S10). In the following, we thus restrict our focus to $\dim H=2$.
@@
We noticed that there were notable differences in the radius of latent space oscillations from colony to colony, two of which are highlighted in (Fig. 7c). To understand this difference, we examined the original microscopy time series corresponding to both large and small latent radius (Fig. 7e-f) and found that while the large-radius sample showed clear, nearly uniform oscillations in intensity, the small-radius samples appeared much more heterogeneous.
@@
To quantify this we consider each pixel to be an independent oscillator, akin to a spatial Kuramoto model [81, 82, 83], and compute a global synchronization order parameter $r(t)$ (see SI Section 7). For each colony we calculate the time-averaged synchronization $\langle r(t)\rangle_{t}$ and find that two clusters emerge corresponding to high and low synchronization (Fig. 7e). These clusters are precisely those representing trajectories of large and small latent radius, suggesting that variational IB learns to encode the synchronization of the colony in the latent variable radius. As a check, we perform IB on a simulated locally-coupled Kuramoto model as a system sharing many features of the experimental system. Here we also learn an encoding in which the latent radius corresponds to the synchronization order parameter (SI Fig. S11).
@@
In the SI we compare the performance of variational IB to several other model reduction methods and find that IB delivers more interpretable and well-behaved features. This is likely due to the fact that many standard methods for data-driven model reduction rely on assumptions about the dynamics which may not be appropriate in the case at hand, such as linearity. Even among deep learning methods free of such assumptions, such as time-lagged autoencoders, the variables learned by IB appear more interpretable. This increased interpretability is likely due to the compression term which effectively regularizes the latent space by encouraging the network to learn slow transfer operator eigenfunctions. While there are many specific variants of DMD [17, 84, 85, 86, 87] or autoencoders for dynamics [28, 29, 30, 88] that may outperform variational IB in some cases, we find that in this real-world example it yields the smoothest and most interpretable latent variables without a tailored pre-processing (SI Fig. S17).
@@ VIII Conclusion
We have related information-theoretical properties of dynamical systems to the spectrum of the transfer operator. We illustrate our findings on several simple and analytically tractable systems, and turn them into a practical tool using variational IB, which learns an encoding variable with a neural network. The latent variables of these networks can be interpreted as transfer operator eigenfunctions even though the network was not explicitly constructed to learn these: it optimizes a purely information-theoretic objective that contains no knowledge of a transfer operator or dynamics. This allows one to harness the power of neural networks to learn physically-relevant latent variables and, combined with methods such as SINDY, even complete reduced models with their governing equations. Biological systems are an ideal setting for such methods: despite their apparent complexity, they can often be captured by low-dimensional descriptions which are difficult to identify by physical considerations alone [3, 62, 89, 90]. We have shown that variational IB is a potentially powerful tool for these cases, and can discover slow variables even directly from image data without significant preprocessing. In the SI we showcase further applications, including to chaotic systems, where they may yield computationally efficient methods to calculate e.g. fractal dimensions.
IX Methods
Information processing inequality
Here we show how to obtain (2) from (1). The variables we are considering form the Markov chain
$X_{t+\Delta t}-X_{t}-h_{\text{enc}}(X_{t})-\mathcal{S}^{\Delta t}h_{\text{enc}}(X_{t})-h^{-1}_{\text{dec}}(\mathcal{S}^{\Delta t}h_{\text{enc}}(X_{t}))$
where $X\,\,\text{---}\,\,Y\,\,\text{---}\,\,Z$ means that $Z$ is conditionally independent of $X$, given $Y$: $p(z,x|y)=p(z|y)p(x|y)$. As a consequence of the data processing inequality, for any transformation $Z=g(Y)$ one has
$I(X,g(Y))\leq I(X,Y).$
where equality is obtained if $g$ is a bijection [58]. Consequently, we have
$I(h^{-1}_{\text{dec}}(\mathcal{S}^{\Delta t}h_{\text{enc}}(X_{t})),X_{t+\Delta t})$ $\leq I(\mathcal{S}^{\Delta t}h_{\text{enc}}(X_{t}),X_{t+\Delta t})$
$\leq I(h_{\text{enc}}(X_{t}),X_{t+\Delta t}),$
where the inequalities are maximized if $h^{-1}_{\text{dec}}$ and $\mathcal{S}^{\Delta t}$ are both bijective functions. Notably, $\mathcal{S}^{\Delta t}=\text{identity}$ is a solution. However, this should not be taken to mean that $H_{t}$ has no dynamics, as it will still have dynamics induced by the evolution of $X_{t}$.
@@ IX Methods
[…]
@@ Acknowledgements
Supplementary Information
I Background material
Mutual information and entropy
$\mathcal{H}(X)=-\int\text{d}x\,p(x)\log p(x).$
High entropy means high uncertainty. The conditional entropy $\mathcal{H}(X|Y)$ between two variables is the average entropy of $p(x|y)$, where the average is taken over $p(y)$:
$\mathcal{H}(X|Y)=-\int\text{d}y\,p(y)\int\text{d}x\,p(x|y)\log p(x|y).$
The mutual information between $X$ and $Y$ corresponds to the (average) reduction of uncertainty about $X$, given knowledge of $Y$:
$I(X,Y)$ $=\mathcal{H}(X)-\mathcal{H}(X|Y)$
$=D_{\text{KL}}(p(x,y)\lVert p(x)p(y))$
$=\int\text{d}x\text{d}y\,p(x,y)\log\frac{p(x,y)}{p(x)p(y)}.$ (S1)
The three formulations of the mutual information above are equivalent and useful in diffent settings. For example, formulation as a Kullback-Leibler divergence immediately tells us that mutual information is non-negative.
Information bottleneck
@@ I Background material
$\mathcal{L}_{\text{IB}}[p(h|x)]$ $=I(X,H)-\beta I(Y,H).$ (S2)
@@
$\begin{split}\mathcal{L}_{\text{IB}}[p(h|x)]&=\sum_{x,h}p(h|x)p(x)\log\frac{p(h|x)}{p(h)}-\beta\sum_{y,h}p(y,h)\log\frac{p(y,h)}{p(h)p(y)}+\sum_{x}\lambda(x)\left(1-\sum_{h}p(h|x)\right).\end{split}$ (S3)
@@
$p(h|x)=\frac{p(h)}{N(x)}\exp\left[\beta\int dy\,p(y|x)\log p(y|h)\right].$ (S5)
When $\beta<1$, it follows from the data processing inequality $I(X,H)\geq I(Y,H)$ that (S2) is minimized by a trivial encoder $p(h|x)=p(h)$. In this case, $\mathcal{L}_{\text{IB}}=0$ and no information passes through the bottleneck. As $\beta$ is increased, more information is allowed through the bottleneck until the relevant variables begin to gain a dependence on the state $x$. This occurs suddenly for a certain value of $\beta=\beta_{1}>1$ at which the encoder becomes non-uniform and $I(H,X)$ becomes non-zero. This is referred to as an IB transition; for increasing $\beta$, there may be a sequence of transitions at $\beta_{2},\beta_{3},$... etc.
Transfer operators
The transfer operator of a Markovian dynamical system describes the evolution of probability distributions
$p_{X_{t+\Delta t}}(x)=U^{\Delta t}[p_{X_{t}}](x)=\int dx^{\prime}p(X_{t+\Delta t}=x|X_{t}=x^{\prime})p(X_{t}=x^{\prime}).$ (S6)
The operator may be decomposed as
@@
For noisy dynamical systems the operator $U^{\Delta t}$ is nearly compact, meaning that its essential spectral radius will be small [49, 50, 33]. We therefore neglect it in the following.
[…]
and they satisfy the biorthogonality condition
$\langle\phi_{n},\rho_{m}\rangle=\delta_{nm}.$
One may use these functions to decompose the conditional probability distribution as
$p(x_{t+\Delta t}|x_{t})=\sum_{n}e^{\lambda_{n}\Delta t}\rho_{n}(x_{t+\Delta t})\phi_{n}(x_{t}).$ (S9)
@@
$\dot{\mathbf{x}}=\mathbf{f}(\mathbf{x})+\sqrt{2D(\mathbf{x})}\mathbf{\xi}(t)$ (S10)
@@
This framework captures purely deterministic dynamics, which are obtained by setting $D=0$. For dynamics given by a Langevin equation, probability distributions evolve according to the corresponding Fokker-Planck equation
@@
Here we recognize $\mathcal{L}$ as the infinitesimal generator of the transfer operator $U$: evolution for a time $t$ is given by $U=\text{e}^{t\mathcal{L}}$. The adjoint operator is given by the so-called backward Kolmogorov equation, and it describes the evolution of functions $\phi$.
$\mathcal{L}^{\dagger}\phi(x)=f_{i}(\mathbf{x})\partial_{i}(\phi(\mathbf{x}))+D(\mathbf{x})C_{ij}\partial_{i}\partial_{j}(\phi(\mathbf{x})).$ (S11)